the sides of a right triangle are a, a+d, a+2d with a and d both positive. what is the value of ratio a:d?

Grade:11

2 Answers

Vikas TU
11 years ago

a, a+d, a+2d are in A.P

but in this qstn. it doesn''t matter.

Just apply the simple and sober Pythogorus formula on these sides of right triangle.

a2 + (a+d)2 = (a + 2d)2

a a2 + d2  +2ad = a+ 4d2 + 4ad

           a2 + d- 2ad = 4d

                    (a - d)4d2

                     a - d =  +-2d

Case 1.    a - d = 2d

                  a = 3d

                  a/d = 3

Case 2.   a-d = -2d

                a= - d

                  or

               d = -a

but given that d > 0

thus, answer is a/d = 3

 

 

Plz approve!!!!!!!!!!!

jaymin prajapati
10 years ago

a^2-2ad+d^2=4^2.so a-d holl squre is eqal to +- 2d.but d>0. so a/d=3.